OpenStudy (anonymous):

The slope of the line whose equation is 3y + 2x = 1 is -2/3 -3/2 1/3

OpenStudy (anonymous):

using the equation y=mx+b we can set the following equation like that to find m the slope and we get -2/3

OpenStudy (anonymous):

3y+2x=1 We need to put your equation in y=mx+b form. -2x -2x _________________ 3y=-2x+1 divide 3 on both sides y=-2x/3+1/3 Can you figure the slope now?